# Normalized Remainders: Definition, Origins, Backgrounds, Recent Results, and Basic Properties

### v1

@@ 1. Definition of the normalized remainder
Let $f$ be an infinitely differentiable real-valued function defined on an interval $I\subseteq\mathbb{R}$ with $x_{0}\in I$ being an inner point. The informal series expansion of $f$ around the point $x_{0}$ is
@@
We call the quantity $T_{n}[f(x),x_{0}]$ the $n$th normalized remainder of the informal series expansion (1.1) of $f$ around $x_{0}$.
@@
The terminology of the normalized remainder $T_{n}[f(x)]$ was first explicitly appeared in the papers [5, Section 5], [33, Section 5], [30, Remark 7], and [41, 43, 47, 70, 72, 75, 76, 81]. It is also known as normalized tail in [30, Remark 7] and [5, 41, 47, 70, 76, 81] and as Qi’s normalized remainder in the papers [43, 75].
@@ 2. Origins in mathematical analysis
We now start out to retrospect origins of inventing the concept of the normalized remainders.
@@ 2.1. First origin related to tangent
This guess was solved and applied in [54] and in the proof of [67, Theorem 3.1].
@@
in [16, p. 55], where $B_{2k}$ denotes the Bernoulli numbers which are generated in [68, p. 3] by
@@ 2.2. Second origin related to sine
In the paper [35, Theorem 1], motivated by the Maclaurin power series expansion of $\ln T_{2}[\tan x]$ obtained in [32], we expanded the logarithm $\ln T_{2}[\sin x]$ into a Maclaurin series. This result can be regarded as a generalization of the series expansion
@@
In the paper [35, Theorem 2], stimulated by the decreasing monotonicity obtained in [54], we proved that the ratio $R_{1}(x)$ defined by (2.4) decreasingly maps from $0,\pi]$ onto $\bigl[0,\frac{3}{10}\bigr]$.
@@ 2.3. Third origin related to cosine
In [33, Theorem 1], we expanded the logarithm $\ln T_{1}[\cos x]$ into a Maclaurin series. This expansion can be regarded as a generalization of the series expansion
@@
In [33, Theorem 2], we proved that the ratio $R_{2}(x)$ defined by (2.7) decreasingly maps $\bigl[0,\frac{\pi}{2}\bigr]$ onto $\bigl[0,\frac{1}{6}\bigr]$.
In [33, Remark 2], Theorem 1 in [33] was simplified as
@@
In [33, Remark 3], we constructed a positive and even function
@@
Further, in the conclusion section [33, Section 5], we first explicitly invented the concept and used the terminology of the normalized remainder as follows.
@@ 3.1. Normalized remainders of sine and cosine functions
In [70, Theorem 1], we expanded the logarithm $\ln T_{2n-1}[\cos x]$ for $n\in\mathbb{N}_{0}$ into a Maclaurin series. This gives a generalization of [33, Theorem 1] and the series expansions (2.8) and (2.9).
In [70, Theorem 2], we proved the following conclusions:
  1. (1)
@@
  2. (2)
    The normalized remainder $T_{1}[\cos x]$ is decreasing on the intervals $(0,2\pi)$ and $(x_{k},2(k+1)\pi)$ for $k\in\mathbb{N}$, while it is increasing on $(2k\pi,x_{k})$ for $k\in\mathbb{N}$, where $x_{k}\in(2k\pi,2(k+1)\pi)$ for $k\in\mathbb{N}$ is the zero of the equation $\tan\frac{x}{2}=\frac{x}{2}$ on $(0,\infty)$.
In [70, Theorem 3], we verified that the ratio
@@
is decreasing on $\bigl[0,\frac{\pi}{2}\bigr]$. This monotonicity result generalizes [33, Theorem 2], because the ratio $R_{3}(x)$ is a generalization of $R_{2}(x)$ defined in (2.7).
@@
  2. (2)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n}[\sin x]$ is positive and decreasing in $x\in(0,\infty)$, while it is concave in $x\in(0,\pi)$.
  3. (3)
    The normalized remainder $T_{1}[\cos x]$ is nonnegative on $(0,\infty)$, is decreasing on $0,2\pi]$, and is concave on $(0,x_{0})$, where $x_{0}\in\bigl(\frac{\pi}{2},\pi\bigr)$ is the first positive zero of the equation
@@
    For $n\geq 2$, the normalized remainder $T_{2n-1}[\cos x]$ is positive and decreasing on $(0,\infty)$ and is concave on $(0,\pi)$.
  4. (4)
@@
  6. (6)
    Using the formulas (3.1) and (3.2), as done in [30, Remark 7], we reformulated in [41, Section 5] all the above results related to the normalized remainders $T_{2n}[\sin x]$ and $T_{2n-1}[\cos x]$ in terms of the generalized hypergeometric function $_{1}F_{2}(a;b,c;z)$.
@@ 3.2. Normalized remainders of tangent function and its square
Recall from the monographs [39, Chapter XIII], [64], and [73, Chapter IV] that,
  1. (1)
@@
A function $f(x)$ is completely monotonic on $(a,b)$ if and only if it is absolutely monotonic on (b,a)(-b,-a); see [73, p. 145, Definition 2c].
In [19, Definition 1] and [55, Definition 1], the concepts of logarithmically absolutely (completely) monotonic functions were defined as follows:
  1. (1)
@@
In [19, Theorem 1], the authors proved that a logarithmically absolutely monotonic function on an interval $I$ is also absolutely monotonic on $I$, but not conversely. In the paper [7], [19, Theorem 4], and [55, Theorem 1], the authors proved that a logarithmically completely monotonic function on an interval $I$ is also completely monotonic on $I$, but not conversely.
@@
  1. (1)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n-2}[\tan z]$ is even on $\bigl(-\frac{\pi}{2},\frac{\pi}{2}\bigr)$, is absolutely monotonic in $z\in\bigl(0,\frac{\pi}{2}\bigr)$, and is completely monotonic in $z\in\bigl(-\frac{\pi}{2},0\bigr)$.
  2. (2)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n-2}[\tan x]$ is logarithmically absolutely monotonic in $x\in\bigl(0,\frac{\pi}{2}\bigr)$.
  3. (3)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n-2}[\tan x]$ is logarithmically completely monotonic in $x\in\bigl(-\frac{\pi}{2},0\bigr)$.
  4. (4)
    For $n\in\mathbb{N}$, the ratio $\frac{(T_{2n-2}[\tan x])^{\prime}}{xT_{2n-2}[\tan x]}$ is absolutely monotonic in $x\in\bigl(0,\frac{\pi}{2}\bigr)$.
  5. (5)
    For $n\in\mathbb{N}$, the ratio $\frac{(T_{2n-2}[\tan x])^{\prime}}{xT_{2n-2}[\tan x]}$ is completely monotonic in $x\in\bigl(-\frac{\pi}{2},0\bigr)$.
  6. (6)
    For $n\in\mathbb{N}$, the function $1-\frac{T_{2n-2}[\tan x]}{T_{2n}[\tan x]}$ is absolutely monotonic in $x\in\bigl(0,\frac{\pi}{2}\bigr)$ and is completely monotonic in $x\in\bigl(-\frac{\pi}{2},0\bigr)$. Consequently, the ratio $\frac{T_{2n-2}[\tan x]}{T_{2n}[\tan x]}$ for $n\in\mathbb{N}$ is decreasing in $x\in\bigl(0,\frac{\pi}{2}\bigr)$, is increasing in $x\in\bigl(-\frac{\pi}{2},0\bigr)$, and is concave in $x\in\bigl(-\frac{\pi}{2},\frac{\pi}{2}\bigr)$.
  7. (7)
@@
In [36], the ratio $\frac{T_{2n+1}[\tan^{2}x]}{T_{2n-1}[\tan^{2}x]}$ for $n\in\mathbb{N}$ was proved to be increasing in $x\in\bigl(-\frac{\pi}{2},0\bigr)$ and to be decreasing in $x\in\bigl(0,\frac{\pi}{2}\bigr)$.
@@ 3.3. Normalized remainders of exponential function
  5. (5)
    For $n\in\mathbb{N}$, the ratio $\frac{B_{2n-1}(t)}{B_{2n+1}(t)}$ is increasing in $t\in\bigl(0,\frac{1}{2}\bigr)$ and decreasing in $t\in\bigl(\frac{1}{2},1\bigr)$, where the classical Bernoulli polynomials $B_{k}(t)$ are generated in [68, p. 3] by
@@
By [20, Theorem 2.1], we deduce that the logarithm $\ln T_{0}[\operatorname{e}^{x}]$ is convex on $(-\infty,\infty)$, is $3$-convex (that is, $\frac{\operatorname{d}^{3}\ln T_{0}[\operatorname{e}^{x}]}{\operatorname{d\!}x^{3}}\geq 0$) on $(-\infty,0)$, and is $3$-concave (that is, $\frac{\operatorname{d}^{3}\ln T_{0}[\operatorname{e}^{x}]}{\operatorname{d\!}x^{3}}\leq 0$) on $(0,\infty)$.
Motivated by [77, Proposition 1], Yang–Qi [74] and Pinelis [44] further investigated the monotonicity and other properties of the ratios
@@
Since 2023, when studying properties of normalized remainders of the Maclaurin power series expansions, we discovered some monotonicity results and inequalities on sequences and functions involving the ratio $\bigl|\frac{B_{2k+2}}{B_{2k}}\bigr|$ and the Riemann zeta function $\zeta(t)$. For more details, please refer to the papers [6] and [50, Section 1.12].
@@
    In [61, Theorem 2.1], a closed-form expression for $A_{j}$ was given by
@@
    $D_{\ell+1}(n)$ $=\begin{vmatrix}\frac{1}{\binom{n+1}{1}}&\frac{\binom{0}{0}}{\binom{n}{0}}&0&0&0&\dotsm&0&0\\ \frac{1}{\binom{n+2}{2}}&\frac{\binom{1}{0}}{\binom{n+1}{1}}&\frac{\binom{1}{1}}{\binom{n}{0}}&0&0&\dotsm&0&0\\ \frac{1}{\binom{n+3}{3}}&\frac{\binom{2}{0}}{\binom{n+2}{2}}&\frac{\binom{2}{1}}{\binom{n+1}{1}}&\frac{\binom{2}{2}}{\binom{n}{0}}&0&\dotsm&0&0\\ \frac{1}{\binom{n+4}{4}}&\frac{\binom{3}{0}}{\binom{n+3}{3}}&\frac{\binom{3}{1}}{\binom{n+2}{2}}&\frac{\binom{3}{2}}{\binom{n+1}{1}}&\frac{\binom{3}{3}}{\binom{n}{0}}&\dotsm&0&0\\ \frac{1}{\binom{n+5}{5}}&\frac{\binom{4}{0}}{\binom{n+4}{4}}&\frac{\binom{4}{1}}{\binom{n+3}{3}}&\frac{\binom{4}{2}}{\binom{n+2}{2}}&\frac{\binom{4}{3}}{\binom{n+1}{1}}&\dotsm&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ \frac{1}{\binom{n+\ell-1}{\ell-1}}&\frac{\binom{\ell-2}{0}}{\binom{n+\ell-2}{\ell-2}}&\frac{\binom{\ell-2}{1}}{\binom{n+\ell-3}{\ell-3}}&\frac{\binom{\ell-2}{2}}{\binom{n+\ell-4}{\ell-4}}&\frac{\binom{\ell-2}{3}}{\binom{n+\ell-5}{\ell-5}}&\dotsm&\frac{\binom{\ell-2}{\ell-2}}{\binom{n}{0}}&0\\ \frac{1}{\binom{n+\ell}{\ell}}&\frac{\binom{\ell-1}{0}}{\binom{n+\ell-1}{\ell-1}}&\frac{\binom{\ell-1}{1}}{\binom{n+\ell-2}{\ell-2}}&\frac{\binom{\ell-1}{2}}{\binom{n+\ell-3}{\ell-3}}&\frac{\binom{\ell-1}{3}}{\binom{n+\ell-4}{\ell-4}}&\dotsm&\frac{\binom{\ell-1}{\ell-2}}{\binom{n+1}{1}}&\frac{\binom{\ell-1}{\ell-1}}{\binom{n}{0}}\\ \frac{1}{\binom{n+\ell+1}{\ell+1}}&\frac{\binom{\ell}{0}}{\binom{n+\ell}{\ell}}&\frac{\binom{\ell}{1}}{\binom{n+\ell-1}{\ell-1}}&\frac{\binom{\ell}{2}}{\binom{n+\ell-2}{\ell-2}}&\frac{\binom{\ell}{3}}{\binom{n+\ell-3}{\ell-3}}&\dotsm&\frac{\binom{\ell}{\ell-2}}{\binom{n+2}{2}}&\frac{\binom{\ell}{\ell-1}}{\binom{n+1}{1}}\end{vmatrix}$
@@
In [78, Theorem 1], we verified the strict inequality
$\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+2}{n+1}}\frac{x^{j}}{j!}\Biggr]\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{j+n}{n}}\frac{x^{j}}{j!}\Biggr]\\ <\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n+1}}\frac{x^{j}}{j!}\Biggr]\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n}}\frac{x^{j}}{j!}\Biggr]$ (3.12)
for $x\in\mathbb{R}$ and $n\in\mathbb{N}_{0}$. In [78, Theorem 2], by virtue of the inequality (3.12), we proved that the ratio $\frac{T_{n+1}[\operatorname{e}^{x}]}{T_{n}[\operatorname{e}^{x}]}$ for $n\in\mathbb{N}_{0}$ is decreasing in $x\in\mathbb{R}$.
Theorem 2 in [78] extends [47, Theorem 1] from $x\in(0,\infty)$ to $x\in\mathbb{R}$.

### v2

# Normalized Remainders: Definition, Origins of Ideas, Historic Backgrounds, Numerous Recent Results, Some Basic Properties, and Related Guesses and Problems

@@ 1. Definition of the normalized remainder
Let $f$ be an infinitely differentiable real-valued function on an interval $I\subseteq\mathbb{R}$ with $x_{0}\in I$ being an inner point. The informal series expansion of $f$ around the point $x_{0}$ is
@@
We call the function $T_{n}[f(x),x_{0}]$ of $x\in I\subseteq\mathbb{R}$ for $n\in\{-1\}\cup\mathbb{N}_{0}$ the $n$th normalized remainder of the informal series expansion (1.1) of $f$ around $x_{0}$.
@@
The terminology “normalized remainder” first explicitly appeared in the papers [5, Section 5], [31, Section 5], [32, Remark 7], and [43, 45, 49, 73, 75, 78, 79, 84]. Earlier literature also referred to this quantity as the “normalized tail”; see [32, Remark 7] and [5, 43, 49, 73, 79, 84]. In addition, the term “Qi’s normalized remainder” has been used in the papers [45, 78], for example.
The reason the term “normalized” is used is because $T_{n}[f(x_{0}),x_{0}]=1$ for all $n\geq-1$ and is for differentiating the quantity $T_{n}[f(x),x_{0}]$ from the classical remainders such as the Taylor remainder, the Lagrange remainder, the Cauchy remainder, the Schlömilch remainder, and the Rouché remainder collected in [81, Section 6]. The notation $T_{n}[f(x)]$ was designed and applied in [21, 22, 23, 33, 37, 53, 68, 74, 81, 82, 83], for example.
@@ 2. Origins in mathematical analysis
We now start out to retrospect the origins of inventing the concept of the normalized remainders $T_{n}[f(x),x_{0}]$.
@@ 2.1. First origin related to tangent
This guess was solved and applied in [57] and in the proof of [70, Theorem 3.1].
@@
in [16, p. 55], where $B_{2k}$ denotes the Bernoulli numbers which are generated in [71, p. 3] by
@@ 2.2. Second origin related to sine
In the paper [36, Theorem 1], motivated by the Maclaurin power series expansion of $\ln T_{2}[\tan x]$ obtained in [34], we expanded the logarithm $\ln T_{2}[\sin x]$ into a Maclaurin series. This result can be regarded as a generalization of the series expansion
@@
In the paper [36, Theorem 2], stimulated by the decreasing monotonicity obtained in [57], we proved that the ratio $R_{1}(x)$ defined by (2.4) decreasingly maps from $0,\pi]$ onto $\bigl[0,\frac{3}{10}\bigr]$.
@@ 2.3. Third origin related to cosine
In [31, Theorem 1], we expanded the logarithm $\ln T_{1}[\cos x]$ into the Maclaurin series. This expansion can be regarded as a generalization of the series expansion
@@
In [31, Theorem 2], we proved that the ratio $R_{2}(x)$ defined by (2.7) decreasingly maps $\bigl[0,\frac{\pi}{2}\bigr]$ onto $\bigl[0,\frac{1}{6}\bigr]$.
In [31, Remark 2], Theorem 1 in [31] was simplified as
@@
In [31, Remark 3], we constructed a positive and even function
@@
Further, in the conclusion section [31, Section 5], we first explicitly invented the concept and used the terminology of the normalized remainder as follows.
@@ 3.1. Normalized remainders of sine and cosine functions
In [73, Theorem 1], we expanded the logarithm $\ln T_{2n-1}[\cos x]$ for $n\in\mathbb{N}_{0}$ into a Maclaurin series. This gives a generalization of [31, Theorem 1] and the series expansions (2.8) and (2.9).
In [73, Theorem 2], we proved the following conclusions:
  1. 1)
@@
  2. 2)
    The normalized remainder $T_{1}[\cos x]$ is decreasing on the intervals $(0,2\pi)$ and $(x_{k},2(k+1)\pi)$ for $k\in\mathbb{N}$, and increasing on $(2k\pi,x_{k})$ for $k\in\mathbb{N}$, where $x_{k}\in(2k\pi,2(k+1)\pi)$ for $k\in\mathbb{N}$ is the zero of the equation $\tan\frac{x}{2}=\frac{x}{2}$ on $(0,\infty)$.
In [73, Theorem 3], we verified that the ratio
@@
is decreasing on $\bigl[0,\frac{\pi}{2}\bigr]$. This monotonicity result generalizes [31, Theorem 2], because the ratio $R_{3}(x)$ is a generalization of $R_{2}(x)$ defined in (2.7).
@@
  2. 2)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n}[\sin x]$ is positive and decreasing in $x\in(0,\infty)$, and concave in $x\in(0,\pi)$.
  3. 3)
    The normalized remainder $T_{1}[\cos x]$ is nonnegative on $(0,\infty)$, decreasing on $0,2\pi]$, and concave on $(0,x_{0})$, where $x_{0}\in\bigl(\frac{\pi}{2},\pi\bigr)$ is the first positive zero of the equation
@@
    For $n\geq 2$, the normalized remainder $T_{2n-1}[\cos x]$ is positive and decreasing on $(0,\infty)$, and concave on $(0,\pi)$.
  4. 4)
@@
  6. 6)
    Using the formulas (3.1) and (3.2), as done in [32, Remark 7], we reformulated in [43, Section 5] all the above results related to the normalized remainders $T_{2n}[\sin x]$ and $T_{2n-1}[\cos x]$ in terms of the generalized hypergeometric function $_{1}F_{2}(a;b,c;z)$.
@@ 3.2. Normalized remainders of tangent function and its square
Recall from the monographs [41, Chapter XIII], [67], and [76, Chapter IV] that,
  1. 1)
@@
A function $f(x)$ is completely monotonic on $(a,b)$ if and only if it is absolutely monotonic on (b,a)(-b,-a); see [76, p. 145, Definition 2c].
Recall from [19, Definition 1] and [58, Definition 1] that,
  1. 1)
@@
In [19, Theorem 1], the authors proved that a logarithmically absolutely monotonic function on an interval $I$ is also absolutely monotonic on $I$, but not conversely. In the paper [7], [19, Theorem 4], and [58, Theorem 1], the authors proved that a logarithmically completely monotonic function on an interval $I$ is also completely monotonic on $I$, but not conversely.
@@
  1. 1)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n-2}[\tan x]$ is even on $\bigl(-\frac{\pi}{2},\frac{\pi}{2}\bigr)$, absolutely monotonic in $x\in\bigl(0,\frac{\pi}{2}\bigr)$, and completely monotonic in $x\in\bigl(-\frac{\pi}{2},0\bigr)$.
  2. 2)
    For $n\in\mathbb{N}$, the normalized remainder $T_{2n-2}[\tan x]$ is logarithmically absolutely monotonic in $x\in\bigl(0,\frac{\pi}{2}\bigr)$ and logarithmically completely monotonic in $x\in\bigl(-\frac{\pi}{2},0\bigr)$.
  3. 3)
    For $n\in\mathbb{N}$, the ratio $\frac{(T_{2n-2}[\tan x])^{\prime}}{xT_{2n-2}[\tan x]}$ is absolutely monotonic in $x\in\bigl(0,\frac{\pi}{2}\bigr)$ and completely monotonic in $x\in\bigl(-\frac{\pi}{2},0\bigr)$.
  4. 4)
    For $n\in\mathbb{N}$, the function $1-\frac{T_{2n-2}[\tan x]}{T_{2n}[\tan x]}$ is absolutely monotonic in $x\in\bigl(0,\frac{\pi}{2}\bigr)$ and completely monotonic in $x\in\bigl(-\frac{\pi}{2},0\bigr)$. Consequently, the ratio $\frac{T_{2n-2}[\tan x]}{T_{2n}[\tan x]}$ for $n\in\mathbb{N}$ is decreasing in $x\in\bigl(0,\frac{\pi}{2}\bigr)$, increasing in $x\in\bigl(-\frac{\pi}{2},0\bigr)$, and concave in $x\in\bigl(-\frac{\pi}{2},\frac{\pi}{2}\bigr)$.
  5. 5)
@@
In [37], the ratio $\frac{T_{2n+1}[\tan^{2}x]}{T_{2n-1}[\tan^{2}x]}$ for $n\in\mathbb{N}$ was proved to be increasing in $x\in\bigl(-\frac{\pi}{2},0\bigr)$ and decreasing in $x\in\bigl(0,\frac{\pi}{2}\bigr)$.
@@ 3.3. Normalized remainders of exponential function
  5. 5)
    For $n\in\mathbb{N}$, the ratio $\frac{B_{2n-1}(t)}{B_{2n+1}(t)}$ is increasing in $t\in\bigl(0,\frac{1}{2}\bigr)$ and decreasing in $t\in\bigl(\frac{1}{2},1\bigr)$, where the classical Bernoulli polynomials $B_{k}(t)$ are generated in [71, p. 3] by
@@
By [20, Theorem 2.1], we deduce that the logarithm $\ln T_{0}[\operatorname{e}^{x}]$ is convex on $(-\infty,\infty)$, $3$-convex (that is, $\frac{\operatorname{d}^{3}\ln T_{0}[\operatorname{e}^{x}]}{\operatorname{d\!}x^{3}}\geq 0$) on $(-\infty,0)$, and $3$-concave (that is, $\frac{\operatorname{d}^{3}\ln T_{0}[\operatorname{e}^{x}]}{\operatorname{d\!}x^{3}}\leq 0$) on $(0,\infty)$.
Motivated by [80, Proposition 1], Yang–Qi [77] and Pinelis [46] further investigated the monotonicity and other properties of the ratios
@@
Since 2023, when studying properties of normalized remainders of the Maclaurin power series expansions, we discovered some monotonicity results and inequalities on sequences and functions involving the ratio $\bigl|\frac{B_{2k+2}}{B_{2k}}\bigr|$ and the Riemann zeta function $\zeta(t)$. For more details, please refer to the papers [6] and [53, Section 1.12].
@@
    In [64, Theorem 2.1], a closed-form expression for $A_{j}$ was given by
@@
    $D_{\ell+1}(n)$ $=\begin{vmatrix}\frac{1}{\binom{n+1}{1}}&\frac{\binom{0}{0}}{\binom{n}{0}}&0&0&\dotsm&0&0\\ \frac{1}{\binom{n+2}{2}}&\frac{\binom{1}{0}}{\binom{n+1}{1}}&\frac{\binom{1}{1}}{\binom{n}{0}}&0&\dotsm&0&0\\ \frac{1}{\binom{n+3}{3}}&\frac{\binom{2}{0}}{\binom{n+2}{2}}&\frac{\binom{2}{1}}{\binom{n+1}{1}}&\frac{\binom{2}{2}}{\binom{n}{0}}&\dotsm&0&0\\ \frac{1}{\binom{n+4}{4}}&\frac{\binom{3}{0}}{\binom{n+3}{3}}&\frac{\binom{3}{1}}{\binom{n+2}{2}}&\frac{\binom{3}{2}}{\binom{n+1}{1}}&\dotsm&0&0\\ \frac{1}{\binom{n+5}{5}}&\frac{\binom{4}{0}}{\binom{n+4}{4}}&\frac{\binom{4}{1}}{\binom{n+3}{3}}&\frac{\binom{4}{2}}{\binom{n+2}{2}}&\dotsm&0&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ \frac{1}{\binom{n+\ell-1}{\ell-1}}&\frac{\binom{\ell-2}{0}}{\binom{n+\ell-2}{\ell-2}}&\frac{\binom{\ell-2}{1}}{\binom{n+\ell-3}{\ell-3}}&\frac{\binom{\ell-2}{2}}{\binom{n+\ell-4}{\ell-4}}&\dotsm&\frac{\binom{\ell-2}{\ell-2}}{\binom{n}{0}}&0\\ \frac{1}{\binom{n+\ell}{\ell}}&\frac{\binom{\ell-1}{0}}{\binom{n+\ell-1}{\ell-1}}&\frac{\binom{\ell-1}{1}}{\binom{n+\ell-2}{\ell-2}}&\frac{\binom{\ell-1}{2}}{\binom{n+\ell-3}{\ell-3}}&\dotsm&\frac{\binom{\ell-1}{\ell-2}}{\binom{n+1}{1}}&\frac{\binom{\ell-1}{\ell-1}}{\binom{n}{0}}\\ \frac{1}{\binom{n+\ell+1}{\ell+1}}&\frac{\binom{\ell}{0}}{\binom{n+\ell}{\ell}}&\frac{\binom{\ell}{1}}{\binom{n+\ell-1}{\ell-1}}&\frac{\binom{\ell}{2}}{\binom{n+\ell-2}{\ell-2}}&\dotsm&\frac{\binom{\ell}{\ell-2}}{\binom{n+2}{2}}&\frac{\binom{\ell}{\ell-1}}{\binom{n+1}{1}}\end{vmatrix}$
@@
In [81, Theorem 1], we verified the strict inequality
$\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+2}{n+1}}\frac{x^{j}}{j!}\Biggr]\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{j+n}{n}}\frac{x^{j}}{j!}\Biggr]\\ <\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n+1}}\frac{x^{j}}{j!}\Biggr]\Biggl[\sum_{j=0}^{\infty}\frac{1}{\binom{j+n+1}{n}}\frac{x^{j}}{j!}\Biggr]$ (3.12)
for $x\in\mathbb{R}$ and $n\in\mathbb{N}_{0}$. In [81, Theorem 2], by virtue of the inequality (3.12), we proved that the ratio $\frac{T_{n+1}[\operatorname{e}^{x}]}{T_{n}[\operatorname{e}^{x}]}$ for $n\in\mathbb{N}_{0}$ is decreasing in $x\in\mathbb{R}$.
Theorem 2 in [81] extends [49, Theorem 1] from $x\in(0,\infty)$ to $x\in\mathbb{R}$.